Find the value zz of a standard Normal variable that satisfies each of the following conditions.

(a) The point zz with 80% of the observations falling below it
z=z=

(b) The point zz with 90% of the observations falling above it

z=z=

please explain how you came to your solution . If you used the calculator please show the steps and what functions you used. Thank you
Answer #1

Solution:-

a) P(Z < z ) = 0.80

Using standard normal table

z =0.84

b)

P(Z > z ) = 0.90

1- P(z < z) =0.90

P(z < z) = 1-0.90 = 0.10

P(z < -1.28) = 0.10

z = -1.28